OpenStudy (anonymous):

X^2+Y-4=0...WHAT ARE THE INTERCEPTS?? and how do I find them??

OpenStudy (campbell_st):

rewrite the equation \[y =4 - x^2\] y- intercepts are when x = 0 substitute x = 0 into the equation to gain the y-intercept x - intercepts are found by letting y = 0 so you need to solve \[4 - x^2 = 0\] this can be factorised... and the values for x are the intercepts

OpenStudy (anonymous):

thanx

OpenStudy (campbell_st):

hope it helps